OpenStudy (shaik0124):

The average of a set of 18 consecutive integers is 22.5. What is the smallest integer in the set?

OpenStudy (anonymous):

total of these integers=22.5*18=405 let first integer =a then last integer=a+17 then sum of 18 integers \[=\frac{ 18*\left( a+a+17 \right) }{ 2 }=18a+153\] \[[S _{n}=\frac{ n \left( a+l \right) }{ 2 }]\] 18a+153=405 find a